OpenStudy (anonymous):

write the following functions as power series, and calculate their open intervals of convergence \(\frac{\sin(x^{3})}{x}\)

OpenStudy (blockcolder):

\[\sin{x}=\sum_{n=0}^\infty (-1)^{n}\frac{x^{2n+1}}{(2n+1)!}\\ \sin{x^3}=\sum_{n=0}^\infty (-1)^{n}\frac{x^{6n+3}}{(2n+1)!}\\ \frac{\sin{x^3}}{x}=\sum_{n=0}^\infty (-1)^{n}\frac{x^{6n+2}}{(2n+1)!}\] Since the R of sin(x) is infinity, then the R of this is also infinity.

OpenStudy (anonymous):

thank you very much